OpenStudy (cutiekawdfi):

write an equation of a parabola that opens up 1. focus 2.5 units from vertex

OpenStudy (cutiekawdfi):

@KylenCarney

OpenStudy (anonymous):

The equation of a parabola with focus above or below the vertex is (x - h)² = 4p(y - k) where the vertex is (h,k) and p = distance from vertex to to focus, taken as a positive number if the focus is above the vertex and negative if the focus is below the vertex. In this case the vertex = (h,k) = (3,-2) By counting the units from the vertex up to the focus we see that p = 6, and is taken positive since the focus is above the vertex, which also means that the parabola opens upward, So the equation is: (x - h)² = 4p(y - k) (x - 3)² = 4(6)(y - (-2)) (x - 3)² = 24(y + 2) That's all you were asked to find. But you might have been asked to graph it and find the directrix. We draw the focal chord (also called "latus rectum") which has length 4p = 4(6) = 24 and whose midpoint is

OpenStudy (cutiekawdfi):

<3